# Enhancing Screen Time Identification in Children with a Multi-View Vision Language Model and Screen Time Tracker

### v2

@@ 4.4 Text generation and screen type identification
We evaluated the generated text description from Llama with the ground truth using BiLingual Evaluation Understudy (BLEU) [38]. BLEU scores evaluate the similarity of the generated description by Llama and the reference caption. A higher BLEU indicates higher similarity to ground truth. BLEU $N$, where $N$ corresponds to contiguous sequences of $N$ word, indicates an evaluation of quality in a context. In Fig. 4, while value decreases when $N$ increases, comparatively high BLEU scores are obtained from BLEU 1 to BLEU 4. Moreover, our MV-VLM model was able to generated scene description with logic and smoothness.
Figure 5 shows representative scene descriptions in each screen type categories. Key words from each scene description is highlighted in red and categories to the three screen types (i.e., TV, smarpthone, and computer). Our vision language model effectively identified various screen types, their spatial relationships, and action events with objects. For instance, it identified a typical setup in the first row where a TV and a fireplace are typical living room setting. In the second row, it captured a scene of a hand interacting with a laptop, recognizing a person seated in front of the laptop. In the final row, the action of holding a cell phone was precisely detected, along with the accurate identification of a smartphone.
@@ 4.5 Comparison with existing methods
To validate the superiority of MV-VLM over existing methods, we compared the performance of identifying screen types with YOLOv8[34] and single-view vision language model (MiniGPT) [35] in Fig. 4 (b). YOLOv8 is an object detection approach that uses convolution neural network (CNN) as backbone to output the bounding box and screen type. It is currently the state of the art that provides high accuracy in object detection. The performance is around 20% lower than our method in accuracy, indicating that VLM is much powerful than conventional CNN-based approach. As shown in Fig. 5, MV-LVM identified spatial relationship among objects (e.g., objects around the screen) and action events (e.g., holding a phone). The capability of identifying spatial relationship among objects and action events is the key that our approach could outperform conventional CNN-based object detection approach.
@@
MiniGPT took single view image as input and used a pre-trained VLM to generate text followed by the same key word extraction and identification process. The performance of MiniGPT is much worse than proposed method, indicating that it is necessary to retain a VLM model with alignment layer specifically for screen type identification. Importantly, we note a consistently higher of 10% accuracy among all types of screens in Fig. 4 (c). The superiority of our method also demonstrates the need for multi-view processing in vision language model as MiniGPT only process images from single-view. In comparison with other VLM, egocentric image sequences have the advantages of acquiring multi-view images over time for processing. Egocentric images are thus considered as an ideal data source for MV-VLM.
@@
Predicted Classes Positive (screen) 75 10
Negative (non) 28 41
4.6 Ablation study
We conducted an ablation study to investigate the contributions of multi-view and multi-caption elements to detection accuracy. The results are summarized in Table 2. For comparative purposes, we modified two models: 1) we removed the view selection module and directly used a single image as input (denoted by Proposed w/o MV, CLIP); 2) we employed three temporally consecutive images as input instead of using the CLIP framework to select multiple views (denoted by Proposed w/o CLIP). It is important to note that the model without the multi-view functionality differs from the MiniGPT described in the previous section, specifically in its underlying large language model architecture. Both alternative methods showed lower accuracy than the proposed model. Without a sophisticated view selection mechanism, the performance can be comparable (i.e., computer) to or even worse (i.e., smartphone) than that achieved with a single-view approach, particularly for smaller screens whose locations can vary across consecutive images. Overall, the results highlight the importance of both the multi-view structure and the CLIP-based view selection in achieving high identification accuracy for different screen types.
4.7 Binary classification
In behavioral study, it is more important to determine the existence of electronic screens rather than identifying the screen types. To this end, we conducted a binary classification test to evaluate the performance of the MV-VLM in detecting the existence of screens or not. We had 154 groups of multi-view images in the test set. The results, including a confusion matrix, are presented in Table 3. The model achieved an overall accuracy of 75.3%. Notably, we achieved a high sensitivity of 88.2%, indicating a stronger capability in detecting the screen existence with a small fraction of false negative detection.
@@ 4.8 Assessment of comfort on wearable device
Table 4: Assessment of Comfort (All values are out of 10)
@@
Emotion (Is it acceptable for children to wear the sensor?) 8.23
Harm (The sensor does not cause pain or tickling.) 8.36
Movement (The sensor doesn’t restrict children’s moving.) 9.23
Anxiety (Children feel secure wearing the device.) 8.2
@@ 5 Discussion
To ensure a fair evaluation on the algorithm development, we exclude samples with severe motion blurring and images accidentally covered by clothes or other objects. We consider those images as outliers which do not correspond to features captured by wearable sensors. In the integration of a hardware system, we plan to use accelerometers to inform the exclusion of low quality blurry images and use ambient light sensors to instruct the system that there are occlusion and the multi-view images are not needed. Future study will seek to integrate behavioral research to measurement obtained from MV-VLM. By correlating our model’s outputs with behavioral data, we aim to deepen our understanding of the interactions between environmental contexts and individual behaviors. This approach is promising not only to validate MV-VLM but also to extend to more complex, behaviorally-driven studies.
@@ 6 Conclusion
We proposed a lightweight and comfortable wearable sensor solution using screen time tracker to monitor children’s screen exposure. We devised a multi-view vision language model to identify existence of screens from egocentric image sequences. We collected data from children’s free living activities. We explored multi-view images to enhance features of screen identification. This multi-view model is integrated with vision language model, generating image descriptions that are related to screen existence. Our experiments indicate superiority in comparison with conventional vision language model and object detection model. Future work will include accumulative measurements of screen time exposure and associate the screen measurements with other factors in behavioral studies.
7 Acknowledgment

### v3

@@ 4.4 Text generation and screen type identification
We evaluated the generated text description from Llama with the ground truth using BiLingual Evaluation Understudy (BLEU) [44]. BLEU scores evaluate the similarity of the generated description by Llama and the reference caption. A higher BLEU indicates higher similarity to ground truth. BLEU $N$, where $N$ corresponds to contiguous sequences of $N$ word, indicates an evaluation of quality in a context. Moreover, our MV-VLM model was able to generated scene description with logic and smoothness.
Figure 5 shows representative scene descriptions in each screen type categories. Key words from each scene description is highlighted in red and categories to the three screen types (i.e., TV, smarpthone, and computer). ChatGPT4 is utlized to smooth the description for scene understanding. Our vision language model effectively identified various screen types, their spatial relationships, and action events with objects. For instance, in the first row, our model identified a laptop when the laptop screen is partially occluded. In this group, the light among the images shows high variance where the first two images are bright while the third image is dark. In the second row, our proposed method identified a typical TV in a living room, even the screen is contorted through a plastic bottle. In the final row, when images are captured during motion, our proposed MV-VLM distinguished the smartphone from the environment in the blurry images.
@@ 4.5 Comparison with existing methods
To validate the superiority of MV-VLM over existing methods, we compared the performance of identifying screen types with a transformer based model (Multi-View Swin Transformer) [45], a convolution neural network (CNN) based model (YOLOv11) [46], a multi-modal VLM (LLaVa) [47], and a modification of CLIP for classification task. The results are reported in Fig. 6. The Multi-View Swin Transformer (MV-Swin-T) is a swin-transformer based algorithm that takes multiple images as input and classifies the input images. YOLOv11 is an object detection approach that uses CNN as backbone to output the bounding box and screen type. It is currently the state of the art that provides high accuracy in object detection. We concatenated the Multi-View images for the YOLOv11 as inputs. LLaVa is a multi-modal model that takes image and text inputs for caption generation. CLIP for Classification (CLIP-fC) is a modification of CLIP by adding a fully-connected layer on top of CLIP features, enabling it for classification tasks. As shown in Fig. 6, our MV-VLM outperformed the existing state-of-the-art (SOAT) algorithms. Moreover, as shown in Fig. 5, MV-LVM identified spatial relationship among objects (e.g., upon a wooden desk) and action events (e.g., holding a smartphone).
@@
LLaVa took image and text as input and generate captions followed by the same key word extraction and identification process. The performance of LLaVa is much worse (around 15%) than proposed method, indicating that it is necessary to retrain a VLM model with alignment layer specifically for screen type identification. The superiority of our method also demonstrates the need for Multi-View processing in VLMs as LLaVa only processes images from single-view. Moreover, as shown in the text generation section in Fig. 6, our model shows better in terms of BLEU scores compared to the LLaVA, which indicates the impotence of finetuning of VLM in the task of screen identification. The BLEU scores are only available for the VLM models such as LLaVA and MV-VLM.
The performance of MV-Swin-T and YOLOv11 are much worse (around 28% and 25%) compared to our algorithm in terms of average accuracy. These two models lack the language model components, which is necessary in this task to achieve optimal performance. CLIP-fC unitizes a pretrained language model and the information from Multi-View images through major voting to acquire the classification results. It outperforms Multi-View Swin Transformer, YOLOv11, and LLaVa, but has worse average accuracy performance (around 9%) compared to our method. The VLM-based models (MV-LVM, LLaVA, and CLIP-fC) out performs non-VLM models, which echoes the necessity of using VLMs for this application.
4.6 Ablation study
We conducted an ablation study to investigate the contributions of components in model architecture and view selection strategy. The results are reported in Table 2. To validate our model design, we separately removed four components from our design, which are LLM, Multi-View inputs, MiniLM, and Swin-Transformer. To remove the LLM in our framework, we employed a fully-connected layer to perform classification of screens based on the visual and text embeddings. Without Multi-View input, we use a single image and its corresponding environment description for text generation and screen type identification. To remove the MiniLM module, we wrapped the environment descriptions and visual embeddings to prompt and directly sent them to the alignment layer. To validate Swin-Transformer in our framework, we used ResNet for extracting the visual features. The results are reported in Model Archicture Ablation in Table 2. As shown in the table, our design outperforms other ablation variants, which echoes the necessity of the LLM, Multi-View image inputs, and Swin-Transformer in the framework.
Table 2: Performance comparison of model architecture ablation and view selection strategy ablation variants across different screen types. The best performance is highlighted by red and the second best is highlighted by blue
Ablation Variant TV Smartphone Laptop Non-screen Acc.
Ours 0.9013 0.8753 0.9240 0.9561 0.9300
Model Architecture Ablations
w/o LLM (use fully-connected layer) 0.8944 0.8660 0.8948 0.8417 0.8650
w/o Multi-View (use single image) 0.8651 0.8813 0.9198 0.8972 0.8917
w/o MiniLM (use direct input text) 0.8614 0.8521 0.9301 0.9218 0.8967
w/o Swin-Transformer (use ResNet) 0.7039 0.8063 0.8615 0.7077 0.7433
View Selection Strategy Ablations
Remove CLIP 0.8647 0.8515 0.8746 0.9525 0.9083
Replace CLIP with ResNet 0.8661 0.8435 0.8777 0.9447 0.9033
Replace CLIP with KMeans 0.8857 0.8470 0.9027 0.9299 0.9050
Moreover, we performed the ablation study on view selection strategies. Without the CLIP for view selection, we explored the following alternative options: 1) used temporally consecutive frames, 2) replaced CLIP with ResNet in view selection, and 3) replaced CLIP with KMeans in view selection. The results are reported in View Selection Strategy section in Table 2. As shown in the table, our Multi-View section design using CLIP embeddings achieves optimal performance.
4.7 Hyperparameter search of $\tau_{h}$ and $\tau_{l}$
In our view selection algorithm, the choice of upper bound $\tau_{h}$ and lower bound $\tau_{l}$ determine the outcome of Multi-View frames. We performed a parameter search to identify the $\tau_{h}$ and $\tau_{l}$. As shown in Fig. 7, the choice of $\tau_{h}$ as 70% and $\tau_{l}$ as 40% leads to the best F1 value and the second best Average Accuracy over different types of screens. It is worth mentioning that our view selection algorithm demonstrates robustness over the $\tau_{h}$ and $\tau_{l}$. In the nine groups of $\tau_{h}$-$\tau_{l}$ spanning from $30\%$ to $80\%$, the average accuracy scores are higher than 0.91 and the average F1 scores are higher than 0.79.
4.8 Screen vs non-screen classification
In behavioral study, it is more important to determine the existence of electronic screens rather than identifying the screen types. To this end, we conducted a binary classification test to evaluate the performance of the MV-VLM in detecting the existence of screens or not. We had 600 groups of Multi-View images in the test set. The results, including a confusion matrix, are presented in Table 3. The model achieved an overall accuracy of 95.5% and a sensitivity of 93.67%, indicating a stronger capability in detecting the screen existence with a small fraction of false negative detection.
Figure 7: The overall accuracy and F1 scores of hyperparameter search of $\tau_{h}$ and $\tau_{l}$.
@@ 4.8 Screen vs non-screen classification
Actual Classes Positive (screen) 281 19
Negative (non) 8 292
@@ 4.9 Assessment of comfort on wearable device
Table 4: Assessment of Comfortness (All values are out of 10)
@@
Emotion
(Is it acceptable for children to wear the sensor?) 8.23
Harm
(The sensor does not cause pain or tickling.) 8.36
Movement
(The sensor doesn’t restrict children’s moving.) 9.23
Anxiety
(Children feel secure wearing the device.) 8.2
@@ 5 Discussion
To ensure a fair evaluation on the algorithm development, we exclude samples with severe motion blurring and images accidentally covered by clothes or other objects. We consider those images as outliers which do not correspond to features captured by wearable sensors. In the integration of a hardware system, we plan to use accelerometers to inform the exclusion of low quality blurry images and use ambient light sensors to instruct the system that there are occlusion and the Multi-View images are not needed. Future study will seek to integrate behavioral research to measurement obtained from MV-VLM. By correlating our model’s outputs with behavioral data, we aim to deepen our understanding of the interactions between environmental contexts and individual behaviors. This approach is promising not only to validate MV-VLM but also to extend to more complex, behaviorally-driven studies. Moreover, in this study, we focused on identifying single screen types. However, our framework naturally supports the identification of multiple screens within a scene. In the future, we will explore the application of identifying multiple screen types. Moreover, we will consider the proximity of the screen in our framework by leveraging the depth information [48], further improving the performance of the proposed framework in the future.
@@ 6 Conclusion
We proposed a lightweight and comfortable wearable sensor solution using screen time tracker to monitor children’s screen exposure. We devised a Multi-View vision language model to identify existence of screens from egocentric image streams. We collected data from children’s free living activities. We explored Multi-View images to enhance features of screen identification. This Multi-View model is integrated with vision language model, generating image descriptions that are related to screen existence. Our experiments indicate superiority in comparison with conventional vision language model and object detection model. Future work will include accumulative measurements of screen time exposure and associate the screen measurements with other factors in behavioral studies.
Acknowledgments